# Let It Flow: Simultaneous Optimization of 3D Flow and Object Clustering

### v2

@@ IV-E Ablation Study
In Table V, we show metrics when adding one enhancement per time. We observe the benefits of components on the final performance on the Argoverse2 dataset. Without rigidity regularization and clusters a), the method resembles flow only to the nearest neighbors in the target point cloud and, therefore, fails to produce meaningful motions. By including the hard cluster rigidity b), we observe physically consistent motions. However, since we enforce rigidity on over-segmented clusters, the objects still deform. With the addition of soft clusters c), we are able to connect nearby points from different hard segments to rigidity terms and reject outliers, expanding the object rigidity. By guiding the hard clusterization with flows d), we acquire a small boost by connecting lonely clusters (divided by occlusions or separated by LiDAR sampling rate) back to the major rigid body.
@@ IV-F Runtime Benefit of Implementing Soft Clusters as Neighborhoods
We analyze the performance-time trade-off in Figure 5 by acquiring the $EPE$ based on inference time on Argoverse1 dataset while using Neural Prior architecture with proposed Soft Clusters to demonstrate runtime difference between iterative per-cluster rigidity [37] and neighborhood parallelization ($\mathcal{L}_{soft}$). The speed was measured on NVIDIA Tesla V100 GPU on complete point clouds. We compare with optimization-based methods FastNSF [21] and MBNSF [37] with distance transform acceleration proposed by [21] and DBSCAN clustering, and Neural prior [20] without distance transform acceleration [21].
Compared to the MBNSFP [37], our method is approximately 10×10$\times$10 × faster while stopping at the same prediction error. Even though their spatial consistency regularization uses the same algorithm of computing the point-to-point displacements with outlier rejection, our method is significantly faster because we design the clusters on the same-sized neighborhoods, which can be efficiently parallelized on GPUs. In contrast to our method, MBNSFP computes eigenvectors of a matrix with variable-sized clusters [9]. Thus, MBNSFP is looping over $c$ clusters $\{N_{0},N_{1},...,N_{c}\}$ one-by-one, whereas our method allows GPU to parallel the $\mathbb{R}^{N\times k\times k}$ tensor computation for $N$ points in point cloud and $k$ neighbors simultaneously.
Figure 5: Flow estimation error vs. run time on Argoverse1 dataset. Our method in green offers a speed-accuracy trade-off that outperforms the converged baselines on the overall EPE metric.

### v3

@@ IV-E Ablation Study
In Table V, we show metrics when adding one enhancement per time. We observe the benefits of components on the final performance on the Argoverse2 dataset. Without rigidity regularization and clusters a), the method resembles flow only to the nearest neighbors in the target point cloud and, therefore, fails to produce meaningful motions. By including the hard cluster rigidity b), we observe physically consistent motions. However, since we enforce rigidity on over-segmented clusters, the objects still deform. With the addition of soft clusters c), we are able to connect nearby points from different hard segments to rigidity terms and reject outliers, expanding the object rigidity. By guiding the hard clusterization with flows d), we acquire a boost by connecting lonely clusters (divided by occlusions or separated by LiDAR sampling rate) back to the major rigid body. Even though the absolute difference value is small for d), it is non-negligible (2.5% relative improvement on dynamic EPE) to the remaining error.
@@ IV-F Runtime Benefit of Implementing Soft Clusters as Neighborhoods
We analyze the performance-time trade-off in Figure 5 by acquiring the $EPE$ based on inference time on Argoverse dataset while using Neural Prior regularizer with proposed Soft Clusters to demonstrate runtime difference between iterative per-cluster rigidity [35] and neighborhood parallelization ($\mathcal{L}_{soft}$). The speed was measured on NVIDIA Tesla V100 GPU on complete point clouds. We compare with optimization-based methods FastNSF [20] and MBNSF [35] with distance transform acceleration proposed by [20] and DBSCAN clustering, and Neural prior [19] without distance transform acceleration [20].
Compared to the MBNSFP [35], our method is approximately 10×10$\times$10 × faster while stopping at the same prediction error. Even though their spatial consistency regularization uses the same algorithm of computing the point-to-point displacements with outlier rejection, our method is significantly faster because we design the clusters on the same-sized neighborhoods, which can be efficiently parallelized on GPUs. In contrast to our method, MBNSFP [35] computes eigenvectors of a matrix with variable-sized clusters [9]. Thus, MBNSFP [35] is looping over $c$ clusters $\{N_{0},N_{1},...,N_{c}\}$ one-by-one, whereas our method allows GPU to parallel the $\mathbb{R}^{N\times k\times k}$ tensor computation for $N$ points in point cloud and $k$ neighbors simultaneously. Method Without any spatial regularization, i.e. FastNP [20], converges sooner with worse final performance. [Uncaptioned image][20][19][35]
Figure 5: Speed performance comparison between the optimization-based methods with Neural Prior regularizer. We show ratio between the performance and time spent of iterations. The Fast Neural Prior [20] converges fastest and Our proposed loss achieves the best performance. All measured with accelerator proposed in [20].